OpenStudy (anonymous):

if f(X) =2x and g(x) = x-4 what is the value of g(f(3))

OpenStudy (anonymous):

f(x)=2(3)=6 then plug 6 in g(x) =6-4=2

OpenStudy (anonymous):

THANK YOU!

OpenStudy (anonymous):

you'r welcome !!